OpenStudy (anonymous):

Write y = 8^x as a logarithmic function. A. log8x = y B. logx8 = y C. log8y = x D. logy8 = x

OpenStudy (anonymous):

First, take logarithms of both sides: \[y = 8^x\]\[\log y = \log8^x\]Then use the law of logarithms which states that loga^b = bloga to get: \[\log y = x\log8\]Then rearrange to get: \[x = \frac{ \log y }{ \log8 }\]

OpenStudy (anonymous):

Whoops, forgot to tell you that the logarithms are to the base 8, so you can simplify to \[x=\log_{8}y\]So the answer is C.